The indicated function y1(x)

 is a solution of the associated homogeneous equation. Use the method of reduction of order to find a second solution y2(x)

 of the homogeneous equation and a particular solution yp(x)

 of the given nonhomogeneous equation.

y'' + y' = 1;    y1 = 1

y2(x)

yp(x)


1
Expert's answer
2021-07-19T18:50:33-0400

Solution

"y'' + y' = 1"

"{d^2y\over dx^2}+{dy\over dx}=1"

First, reduce the differential equation to first order

Let "z={dy\over dx} \implies {dz\over dx}={d^2y\over dx^2}"

Substituting to differential equation, we have

"{dz\over dx}+z=1"


"{dz\over dx}=1-z"


Do reciprocal on both sides of the differential equation

"{1\over 1-z}={dx\over dz}"

Multiply both sides of the differential equation by "dz" to get

"{dz\over 1-z}=dx"

Integrate both sides

"\int{dz\over 1-z}=\int dx"

"\implies -ln(1-z)=x+C"

"\implies ln(1-z)=-x+C"

Introduce exponential on both sides

"e^{ln(1-z)}=e^{-x+c}"

"\implies 1-z=e^{-x+C} \implies z=1-e^{-x+C}=1-Ce^{-x}"

But "z={dy\over dx}"


"\implies {dy\over dx}=1-Ce^{-x}"

Multiply both sides of the differential equation by "dx"

"\implies dy=(1-Ce^{-x})dx"

Integrate both sides

"\int dy=\int (1-Ce^{-x})dx"

"\implies y=x+C_1e^{-x}+C_2"

"\therefore y_2(x)=e^{-x}" and "y_p(x)=x"